Solve the 3x3 system of equations.
x + y + z = 6
x = 2y
z = x + 1

Answer by TimothyLamb(4379)

x + y + z = 6
x = 2y
z = x + 1
---
put the system of linear equations into standard form
---
x + y + z = 6
x - 2y = 0
x - z = -1
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 2
y = 1
z = 3
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---